OpenStudy (anonymous):

Help with changing equations?? Change the equation -2(x-2)^2+2 into general and standard form of a quadratic equation.

OpenStudy (anonymous):

You first need to FOIL the (x-2)^2 : (x-2)(x-2) = x^2 - 4x +4 Now, multiply the -2 by the result obtained from FOIL-ing -2(x^2 - 4x +4) = -2x^2 + 8x - 8 Now we have: -2x^2 + 8x - 8 + 2 = -2x^2 + 8x -6 Remember the standard form of a quadratic is: ax^2 + bx + c = 0 where a, b, and c are constants. So your answer will be: -2x^2 + 8x + 6 = 0

OpenStudy (anonymous):

My mistake, it's not -2x^2 + 8x + 6 it's -2x^2 + 8x - 6.

OpenStudy (anonymous):

Final answer: -2x^2 + 8x - 6 = 0

OpenStudy (anonymous):

Thank you. So the standard form is -2x^2 + 8x - 6 = 0, then what is the general form?

OpenStudy (anonymous):

No worries! The general form is f(x) = ax^2 + bx + c So yours would be: f(x) = -2x^2 + 8x - 6

OpenStudy (anonymous):

Alright, thank you again!

OpenStudy (anonymous):

You're welcome!